### v1

# Preference-CFR: Beyond Nash Equilibrium for Better Game Strategies

Abstract
Recent advancements in artificial intelligence (AI) have leveraged large-scale games as benchmarks to gauge progress, with AI now frequently outperforming human capabilities. Traditionally, this success has largely relied on solving Nash equilibrium (NE) using variations of the counterfactual regret minimization (CFR) method in games with incomplete information. However, the variety of Nash equilibria has been largely overlooked in previous research, limiting the adaptability of AI to meet diverse human preferences. To address this challenge, where AI is powerful but struggles to meet customization needs, we introduce a novel approach: Preference-CFR, which incorporates two new parameters: preference degree and vulnerability degree. These parameters allow for greater flexibility in AI strategy development without compromising convergence. Our method significantly alters the distribution of final strategies, enabling the creation of customized AI models that better align with individual user needs. Using Texas Hold’em as a case study, our experiments demonstrate how Preference CFR can be adjusted to either emphasize customization, prioritizing user preferences, or to enhance performance, striking a balance between the depth of customization and strategic optimality.
1 Introduction
In machine learning, complex gaming problems are important benchmarks for assessing artificial intelligence (AI). Prominent games such as Chess Hsu (2002), Go Silver et al. (2017, 2016, 2018), StarCraft Vinyals et al. (2019), and Texas Hold’em Moravík et al. (2017); Bowling et al. (2015); Brown and Sandholm (2019b) have significantly influenced both academic research and public interest. Traditionally, research has centered on finding Nash equilibrium (NE), as it guarantees that no player can increase their expected payoff by unilaterally changing strategies. From the perspective of expected payoffs, NE represents the optimal solution in games, leading many studies to regard a game problem as solved once its NE is identified. However, maximizing expected payoffs in the worst case is not the sole criterion for evaluating a strategy’s quality. Here, we introduce two additional indicators beyond expected payoffs.
  1. 1.
    In many games, multiple NE can exist. In economics, exploring the diversity of NE is often more valuable than simply identifying them. For example, Schelling’s work on predicting the emergence of specific NE earned him the 2005 Nobel Prize in Economics. Similarly, the current development of AI demands not only optimal solutions but also diverse and flexible strategies. The goal is to move away from rigid, overly rational AI behaviors and towards strategies that exhibit more human-like characteristics. Moreover, making AI algorithms more interpretable is essential for enhancing their reliability and practical usefulness across various applications.
  2. 2.
    NE focuses solely on the magnitude of expected payoffs, disregarding the variability in those payoffs. However, balancing risk and reward is a crucial aspect of decision-making. In different situations, we may need to adjust this balance and choose strategies accordingly. Relying on NE alone does not provide the flexibility to accommodate varying risk preferences.
Previous algorithms that have achieved success in incomplete information have not addressed the issues mentioned above. To overcome these limitations, we propose a new algorithm called Preference Counterfactual Regret Minimization (Pref-CFR). This algorithm introduces two additional parameters for strategy selection: the preference degree $\delta$, which represents the player’s inclination towards a particular action, and the vulnerability degree $\beta$, which indicates the maximum level of exploitability the player is willing to accept.
More importantly, by setting the preference and vulnerability degrees, Pref-CFR can achieve specific strategy styles as desired by humans (e.g., a highly aggressive play style in poker). Additionally, the implementation of Pref-CFR is straightforward, requiring only minimal code changes from the original CFR, and it remains compatible with many previous CFR variants.
In the experimental section, we first illustrate the limitations of the original CFR in failing to converge to different equilibria. We then highlight the capability of our algorithm to converge to various strategy styles in Texas Hold’em poker.
2 Related Work
The leading algorithm for normal-form games is Regret Minimization (RM), along with its variant, Counterfactual Regret Minimization (CFR) Zinkevich et al. (2007b), which is used in extensive-form games. RM/CFR can converge to NE in two-player zero-sum games and coarse correlated equilibria (CCE) in two/multi-player general-sum games Farina et al. (2021). Notable variants include CFR+ Tammelin (2014), Monte-Carlo CFR (MCCFR) Lanctot et al. (2009), and Discount CFR (DCFR) Brown and Sandholm (2019a). In particular, CFR+ and MCCFR have been instrumental in major AI advancements over the past decade Brown and Sandholm (2019b); Bowling et al. (2015); Brown and Sandholm (2017). In addition to the RM algorithm, the Fictitious Play (FP) algorithm is another commonly used method for solving games. FP was first introduced in Brown’s 1951 paper Berger (2007); Brown (1951), and the treatise The Theory of Learning in Games Fudenberg and Levine (1998) further consolidated prior research, establishing a standardized framework for FP. Recently, Qi et al. combined CFR with the FP algorithm to propose the CFVFP algorithm Qi et al. (2023). Nonetheless, all these methods primarily focus on solving for NE without addressing the applicability and diversity of NE discussed earlier.
In game theory, numerous studies have explored types of equilibria beyond NE. For instance, in 2005, the Nobel Prize in Economics was awarded to Aumann and Schelling for their contributions to enhancing our understanding of game equilibria and improving the concept of NE. Aumann noted that correlated equilibrium can yield fairer and more profitable outcomes compared to NE Aumann (1974). Schelling examined the likelihood of different NE arising in scenarios with multiple equilibria Schelling (1980). Additionally, Fudenberg et al. argued that stable states reached during the learning (or evolutionary) process can also be considered equilibria Fudenberg and Levine (1998). Recently, Ganzfried introduced a new concept known as safe equilibrium, which incorporates the irrationality of opponents and allows for more flexible responses to various adversaries Ganzfried (2023). Nonetheless, while these studies present innovative ideas and perspectives on game problems, they often lack a detailed discussion on how to effectively solve for these new equilibria.
This paper will combine the research of these two aspects to discuss how to converge to different equilibria in theory of learning in games.
3 Notion and Preliminaries
3.1 Game Theory
3.1.1 Normal-Form Game
Normal-form game is the fundamental model in game theory. Let $\mathcal{N}=\{1,2,\dots,i,\dots\}$ denote the set of players, where player $i$ has a finite action set $\mathcal{A}^{i}$. The strategy $\sigma^{i}$ of player $i$ is defined as a $|\mathcal{A}^{i}|$-dimensional probability distribution over $\mathcal{A}^{i}$ (where ||$\cdot$|| ⋅ | represents the number of elements in the set), with $\sigma^{i}(a^{\prime})$ indicating the probability of player $i$ choosing action $a^{\prime}$. The strategy set of player $i$ is denoted as $\Sigma^{i}$, where $\sigma^{i}\in\Sigma^{i}$. Strategies can be categorized into pure strategies and mixed strategies: a pure strategy involves taking a specific action with 100% probability, while all strategies other than pure strategies are considered mixed strategies. A strategy profile $\sigma=\mathop{\times}\limits_{i\in\mathcal{N}}\sigma^{i}$ is a collection of strategies for all players, and $\sigma^{-i}=(\sigma^{1},\dots,\sigma^{i-1},\sigma^{i+1},\dots)$ refers to all strategies in $\sigma$ except for player $i$. The set of all strategy profiles is denoted as $\Sigma=\mathop{\times}\limits_{i\in\mathcal{N}}\Sigma^{i}$. We define the finite payoff function $u^{i}:\Sigma\rightarrow\mathbb{R}$, where $u^{i}(\sigma^{i},\sigma^{-i})$ represents the payoff received by player $i$ when player $i$ selects strategy $\sigma^{i}$ and all other players follow the strategy profile $\sigma^{-i}$. Finally, we define $L=\max_{\sigma\in\Sigma,i\in\mathcal{N}}u^{i}(\sigma)-\min_{\sigma\in\Sigma,i\in\mathcal{N}}u^{i}(\sigma)$ as the payoff interval of the game.
3.1.2 Extensive-Form Games
In extensive-form games, commonly depicted as game trees, the set of players is represented as $\mathcal{N}=\{1,2,\dots\}$. The nodes $s$ within the game tree signify possible states, collectively forming the state set $s\in\mathcal{S}$, while leaf nodes $z\in\mathcal{Z}$ denote terminal states. For each state $s\in\mathcal{S}$, the successor edges delineate the action set $\mathcal{A}(s)$ accessible to either a player or chance events. The player function $P:\mathcal{S}\rightarrow\mathcal{N}\cup\{c\}$ specifies which entity acts at a particular state, where $c$ represents chance.
Information sets $I\in\mathcal{I}^{i}$ comprise collections of states that player $i$ cannot distinguish from one another. The payoff function $R:\mathcal{Z}\rightarrow\mathbb{R}^{|\mathcal{N}|}$ assigns a payoff vector for the players based on the terminal states. The behavioral strategy $\sigma^{i}(I)\in\mathbb{R}^{|\mathcal{A}(I)|}$ is defined as a probability distribution over each information set $I$ for all $I\in\mathcal{I}^{i}$. We also define $\pi_{\sigma}(I)$ as the probability of encountering information set $I$ when all players select actions according to the strategy $\sigma$.
3.1.3 Nash Equilibrium in Games
The best response (BR) strategy for player $i$ in relation to the strategy profile $\sigma^{-i}$ is defined as:
$b^{i}(\sigma^{-i})={\arg\max}_{a^{*}\in\mathcal{A}^{i}}u^{i}(a^{*},\sigma^{-i}).$ (1)
The BR strategy can either be a pure strategy or a mixed strategy; yet, identifying a pure BR strategy is generally more straightforward. For the purposes of our analysis, we will assume that $b(\sigma)$ is a pure strategy. In this context, $\arg\max$ denotes the action that produces the highest payoff within a given set. If there are multiple actions that yield this maximum payoff, the strategy that appears first in lexicographic order will be selected.
The exploitability $\epsilon^{i}$ of player $i$ with respect to a strategy profile $\sigma$ is defined as:
$\epsilon^{i}=u^{i}(b^{i}(\sigma^{-i}),\sigma^{-i})-u^{i}(\sigma),$ (2)
while the overall exploitability $\epsilon$ across all players is calculated as:
$\epsilon=\frac{1}{|\mathcal{N}|}{\sum}_{i\in\mathcal{N}}\epsilon^{i},$ (3)
if $\epsilon=0$, the strategy profile $\sigma$ is a NE; otherwise, it is termed an $\epsilon$-NE. If an algorithm ensures that the exploitability of the game meets the condition $\epsilon\leq CT^{-1}$ after $T$ iterations (where $C$ is a constant), then the convergence rate of this algorithm is $O(T^{-1})$.
3.2 Counterfactual Regret Minimization
We define the counterfactual value $u(I,\sigma)$ as the expected value conditioned on reaching the information set $I$ while all players adopt the strategy $\sigma$, with the exception that player $i$ plays in a way that allows reaching $I$. For every action $a\in\mathcal{A}^{i}(I)$, we denote $\sigma|_{I\rightarrow a}$ as the strategy profile that is identical to $\sigma$ except that player $i$ always selects action $a$ when in information set $I$. The average counterfactual regret defined as:
$\bar{R}_{T}^{i}(I,a)=\frac{1}{T}\sum_{t=1}^{T}\pi_{\sigma_{t}}^{-i}(I)\left(u^{i}\left(I,\sigma_{t}|_{I\rightarrow a}\right)-u^{i}\left(I,\sigma_{t}\right)\right),$ (4)
where $\pi_{\sigma_{t}}^{-i}(I)$ is the probability of information set $I$ occurring given that all players (including chance, except for player $i$) choose actions according to $\sigma_{t}$. We define $\bar{R}_{T}^{i,+}(I,a)=\max(\bar{R}_{T}^{i}(I,a),0)$. The strategy for player $i$ at time $T+1$ is given by:
$\sigma_{T+1}^{i}(I,a)=\begin{cases}\frac{\bar{R}_{T}^{i,+}(I,a)}{\sum_{a\in\mathcal{A}(I)}\bar{R}_{T}^{i,+}(I,a)}&\text{ if }\bar{R}_{T}^{i,+}(I)\neq\textbf{0}\\ \frac{1}{|\mathcal{A}(I)|}&\text{ otherwise, }\end{cases}$ (5)
where $\mathbf{0}$ denotes the zero vector. The average strategy $\bar{\sigma}_{T}^{i}(I)$ for an information set $I$ after $T$ iterations is defined as:
$\bar{\sigma}_{T}^{i}(I)=\frac{\sum_{t=1}^{T}\pi_{\sigma_{t}}^{i}(I)\sigma_{t}^{i}(I)}{\sum_{t=1}^{T}\pi_{\sigma_{t}}^{i}(I)}.$ (6)
Ultimately, as $T\to\infty$, $\bar{\sigma}_{T}$ will converge to a NE.
However, the original CFR method requires traversing the entire game tree, significantly limiting its applicability. In practical scenarios, the Monte Carlo method is often employed, which involves traversing only a subset of the game tree during each iteration. In the MCCFR approach Brown (2020), some players are designated as traversers while others act as samplers during an iteration. The traversers utilize the CFR algorithm to update the regret and average strategy for the information sets they have encountered. Conversely, at the nodes of the samplers and at chance nodes, only a single action is explored, sampled according to the strategy of the sampler player.
3.3 Blackwell Approachability Game
Definition 1
A Blackwell approachability game in normal-form two-player games can be described as a tuple $(\Sigma,u,S^{1},S^{2})$, where $\Sigma$ is a strategy profile, $u$ is the payoff function, and $S^{i}=\mathbb{R}_{\leq 0}^{|\mathcal{A}^{i}|}$ is a closed convex target cone. The Player $i$’s regret vector of the strategy profile $\sigma$ is $R^{i}(\sigma)\in\mathbb{R}^{|\mathcal{A}^{i}|}$, for each component $R^{i}(\sigma)(a_{x})=u^{i}\left(a_{x},\sigma^{-i}\right)-u^{i}\left(\sigma\right)$ , $a_{x}\in\mathcal{A}^{i}$ the average regret vector for players $i$ to take actions at $T$ time $a$ is $\bar{R}_{T}^{i}$
$\bar{R}_{T}^{i}=\frac{1}{T}\sum_{t=1}^{T}R^{i}(\sigma_{t}),$ (7)
at each time $t$, the two players interact in this order:
  • Player 1 chooses a strategy ${\sigma}_{t}^{1}\in\Sigma^{1}$;
  • Player 2 chooses an action ${\sigma}_{t}^{2}\in\Sigma^{2}$ , which can depend adversarially on all the ${\sigma}^{t}$ output so far;
  • Player 1 gets the vector value payoff $R^{1}(\sigma_{t})\in\mathbb{R}^{|\mathcal{A}^{1}|}$.
The goal of Player 1 is to select actions ${\sigma}_{1}^{1},{\sigma}_{2}^{1},\ldots\in\Sigma^{1}$ such that no matter what actions ${\sigma}_{1}^{2},{\sigma}_{2}^{2},\ldots\in\Sigma^{2}$ played by Player 2, the average payoff vector converges to the target set $S^{1}$.
$\min_{\hat{\boldsymbol{s}}\in S^{1}}\left\|\hat{\boldsymbol{s}}-\bar{R}_{T}^{1}\right\|_{2}\rightarrow 0\quad\text{ as }\quad T\rightarrow\infty.$ (8)
Before explaining how to choose the action $\sigma_{t}$ to ensure this goal achieve, we first need to define the forceable half-space:
Definition 2
Let $\mathcal{H}\subseteq\mathbb{R}^{d}$ as half-space, that is, for some $\boldsymbol{a}\in\mathbb{R}^{d}$, $b\in\mathbb{R}$, =$\mathcal{H}=\left\{\boldsymbol{x}\in\mathbb{R}^{d}:\boldsymbol{a}^{\top}\boldsymbol{x}\leq b\right\}$. In Blackwell approachability games, the halfspace $\mathcal{H}$ is said to be forceable if there exists a strategy $\sigma^{i*}\in\Sigma^{i}$ of Player $i$ that guarantees that the regret vector $R^{i}(\sigma)$ is in $\mathcal{H}$ no matter the strategy played by Player $-i$, such that
${R}^{i}\left({\sigma}^{i*},\hat{\sigma}^{-i}\right)\in\mathcal{H}\quad\forall\hat{\sigma}^{-i}\in\Sigma^{-i},$ (9)
and $\sigma^{i*}$ is forcing action for $\mathcal{H}$.
Blackwell’s approachability theorem states the following.
Theorem 1
Goal 8 can be attained if and only if every halfspace $\mathcal{H}_{t}\supseteq S$ is forceable.
The relationship between Blackwell approachability and no-regret learning is:
Theorem 2
Any strategy (algorithm) that achieves Blackwell approachability can be converted into an algorithm that achieves no-regret, and vice versa Abernethy et al. (2011)
If the algorithm achieves Blackwell approachability, the average strategy $\bar{\sigma}_{T}^{i}$ will converge to equilibrium with $T\rightarrow\infty$. The rate of convergence is ${\epsilon_{T}^{i}}\leq\bar{R}_{T}^{i}\leq L\sqrt{|\mathcal{A}^{i}|}/\sqrt{T}$.
4 Method
4.1 Preference CFR
We now introduce the Pref-CFR algorithm. For each information set, we define a preference degree $\delta(I)\in\mathbb{R}^{|\mathcal{A}(I)|}$, where $\delta(I,a)\geq 1$. When $\delta(I,a)=1$ for all $a\in\mathcal{A}(I)$ and $I\in\mathcal{I}$, the algorithm reduces to the original CFR. When $\sum_{a\in\mathcal{A}(I)}\delta(I,a)\neq|\mathcal{A}(I)|$, the strategy for the next iteration $T+1$ is defined as:
$\sigma_{T+1}^{i}(I,a)=\begin{cases}\frac{\delta(I,a)\bar{R}_{T}^{i,+}(I,a)}{\sum_{a\in\mathcal{A}(I)}\delta(I,a)\bar{R}_{T}^{i,+}(I,a)}&\text{ if }\bar{R}_{T}^{i,+}(I)\neq\textbf{0}\\ \frac{\delta(I,a)-1}{\sum_{a\in\mathcal{A}^{i}(I)}{\delta(I,a)-1}}&\text{ otherwise. }\end{cases}$ (10)
Additionally, we can also adopt the best response (BR) strategy for the next iteration:
$\sigma^{i}_{T+1}(I)={\arg\max}_{a\in\mathcal{A}^{i}(I)}\delta(I,a)\bar{R}_{T}^{i}(I,a).$ (11)
In this paper, we denote Pref-CFR(RM) to indicate that the next strategy follows the regret minimization approach, as shown in Equation 10, while Pref-CFR(BR) signifies that the next strategy is based on the BR approach, as presented in Equation 11. In the next section, we will demonstrate that this algorithm can still converge to a NE. In normal-form games, the convergence rate is $O(\delta^{*}T^{-\frac{1}{2}})$, where $\delta^{*}=\max_{a\in\mathcal{A}}\delta(a)$.
If the game has only one equilibrium, or if action $a^{\prime}$ is a dominated strategy, the choice of the preference degree $\delta(I,a^{\prime})$ does not influence the convergence to a NE. Nonetheless, in cases where the game has multiple NEs and the selection probabilities for action $a^{\prime}$ vary across these equilibria, we can deliberately control the final strategy’s direction by adjusting the preference degrees $\delta(I,a^{\prime})$. Specifically, setting a higher preference degree $\delta(I,a^{\prime})$ for a particular action $a^{\prime}$ will increase its selection frequency, thereby raising its probability in the average strategy. Conversely, if the preference degree for this action is lower than that of others, its selection probability will diminish.
While the Pref-CFR algorithm can converge to different NEs in games with multiple equilibria, it will converge to the unique NE if only one exists. To address this, we introduce the vulnerability degree $\beta(I)$, which measures the algorithm’s distance from the NE. We define:
$\bar{B}_{T}^{i}(I,a)=\bar{R}_{T}^{i}(I,a)-\beta(I),$ (12)
the strategy for time $T+1$ is determined as follows:
$\sigma_{T+1}^{i}(I,a)=\begin{cases}{\arg\max}_{a\in\mathcal{A}^{i}(I)}\bar{B}_{T}^{i}(I,a)&\text{ if }\bar{B}_{T}^{i,+}(I,a)\neq{0}\\ \frac{\delta(I,a)-1}{\sum_{a\in\mathcal{A}^{i}(I)}{\delta(I,a)-1}}&\text{ otherwise, }\end{cases}$ (13)
where $\bar{B}_{T}^{i,+}(I,a)=\max\{\bar{B}_{T}^{i}(I,a),0\}$. In the next section, we will prove that in normal-form games, the vulnerability degree satisfies $\epsilon_{\beta}=\max_{I\in\mathcal{I}}\beta(I)$, ensuring that the final strategy is an $\epsilon_{\beta}$-NE strategy. Unlike a NE strategy, which may represent a single point in the simplex, an $\epsilon$-NE strategy must form a convex polyhedron. This provides a wider range of strategic options. By integrating preference and vulnerability, we can devise a strategy that not only aligns with desired stylistic requirements but also minimizes potential losses, even when it is highly targeted.
4.2 Pref-CFR Achieves Blackwell Approachability
Figure 1: The differences in the selected forcing actions and forcing half-spaces of RM, FP and Pref-RM in the two-dimensional plane.
The proof concept for the Pref-CFR algorithm is as follows. According to Paper Zinkevich et al. (2007a), if an algorithm can ensure that regret converges to zero in normal-form games, then applying this algorithm in extensive-form games will lead to the convergence of counterfactual regret to zero (i.e., convergence to a NE). Furthermore, Theorem 2 indicates that any algorithm satisfying Blackwell approachability is equivalent to a no-regret algorithm. Thus, to demonstrate that Pref-CFR converges to NE in extensive-form games, we only need to establish that it satisfies Blackwell approachability in normal-form games. In this proof, we primarily show that Pref-CFR(BR) meets the criteria for Blackwell approachability, while Pref-CFR(RM) can derive similar results through a relatively straightforward conversion.
Let $\bar{R}_{t}^{i,*}$ represent any component of the vector $\bar{R}_{t}^{i}$ such that $\bar{R}_{t}^{i,*}>0$. Next, we identify the point $\boldsymbol{\psi}_{t}=\bar{R}_{t}^{i}-\bar{R}_{t}^{i,*}\in\mathbb{R}^{|\mathcal{A}^{i}|}$ on the axis (noting that this point is not necessarily located on the surface of the target cone $S^{i}$). We can define the normal vector as $\frac{\bar{R}_{t}^{i}-\boldsymbol{\psi}_{t}}{\left|\bar{R}_{t}^{i}-\boldsymbol{\psi}_{t}\right|}$, which allows us to determine the half-space defined by this normal vector and point $\boldsymbol{\psi}_{t}$:
$\mathcal{H}_{t}^{\text{P}}=\left\{\boldsymbol{z}\in\mathbb{R}^{|\mathcal{A}^{-i}|}:(\bar{R}_{t}^{i}-\boldsymbol{\psi}_{t})^{\top}\boldsymbol{z}\leq(\bar{R}_{t}^{i}-\boldsymbol{\psi}_{t})^{\top}\boldsymbol{\psi}_{t}\right\}.$ (14)
Since $\bar{R}_{t}^{i}-\boldsymbol{\psi}_{t}=\bar{R}_{t}^{i,*}$ and $(\bar{R}_{t}^{i}-\boldsymbol{\psi}_{t})^{\top}\boldsymbol{\psi}_{t}=0$, we can simplify this to:
$\mathcal{H}_{t}^{\text{P}}=\left\{\boldsymbol{z}\in\mathbb{R}^{|\mathcal{A}^{-i}|}:\left\langle\bar{R}_{t}^{i,*},\boldsymbol{z}\right\rangle\leq 0\right\},$ (15)
for any point $\boldsymbol{s}^{\prime}\in S^{i}$ there is $\left\langle\bar{R}_{t}^{i,*},\boldsymbol{s}^{\prime}\right\rangle\leq 0$. Then we need to find the forcing action that matches $\mathcal{H}^{\text{P}}_{t}$. According to Definition 2, we need to find a $\sigma^{i*}_{t+1}\in\Sigma^{i}$ that achieves $R^{i}\left(\sigma_{t+1}^{i*},\hat{\sigma}_{t+1}^{-i}\right)\in\mathcal{H}_{t+1}^{i,\text{P}}$ for any $\hat{\sigma}_{t+1}^{-i}\in\Sigma^{-i}$. For simplicity, let $\boldsymbol{\ell}$$=[u^{i}\left(a_{1},\sigma^{-i}\right)$, $\dots]^{\top}\in\mathbb{R}^{|\mathcal{A}^{i}|}$, we rewrite the regret vector as $R^{i}\left(\sigma_{t+1}^{i*},\hat{\sigma}_{t+1}^{-i}\right)=\boldsymbol{\ell}-\left\langle\boldsymbol{\ell},\sigma^{i*}_{t+1}\right\rangle\mathbf{1}$, we are looking for a $\sigma^{i*}_{t+1}\in\Sigma^{i}$ such that:
$R^{i}\left(\sigma_{t+1}^{i*},\hat{\sigma}_{t+1}^{-i}\right)\in\mathcal{H}_{t}^{\text{P}}$ (16)
$\Longleftrightarrow$ $\left\langle\bar{R}_{t}^{i,*},\boldsymbol{\ell}-\left\langle\boldsymbol{\ell},\sigma_{t+1}^{i*}\right\rangle\mathbf{1}\right\rangle\leq 0$
$\Longleftrightarrow$ $\left\langle\bar{R}_{t}^{i,*},\boldsymbol{\ell}\right\rangle-\left\langle\boldsymbol{\ell},\sigma_{t+1}^{i*}\right\rangle\left\langle\bar{R}_{t}^{i,*},\mathbf{1}\right\rangle\leq 0$
$\Longleftrightarrow$ $\left\langle\bar{R}_{t}^{i,*},\boldsymbol{\ell}\right\rangle-\left\langle\boldsymbol{\ell},\sigma_{t+1}^{i*}\right\rangle\left\|\bar{R}_{t}^{i,*}\right\|_{1}\leq 0$
$\Longleftrightarrow$ $\left\langle\boldsymbol{\ell},\frac{\bar{R}_{t}^{i,*}}{\left\|\bar{R}_{t}^{i,*}\right\|_{1}}\right\rangle-\left\langle\boldsymbol{\ell},\sigma_{t+1}^{i*}\right\rangle\leq 0$
$\Longleftrightarrow$ $\left\langle\boldsymbol{\ell},\frac{\left[\bar{R}_{t}^{i}\right]^{*}}{\left\|\left[\bar{R}_{t}^{i}\right]^{*}\right\|_{1}}-\sigma_{t+1}^{i*}\right\rangle\leq 0.$
Therefore, the strategy $\sigma^{i*}_{t+1}=\frac{\bar{R}_{t}^{i,*}}{\left\|\bar{R}_{t}^{i,*}\right\|_{1}}$ can guarantee $\mathcal{H}_{t+1}^{\text{P}}$ to be forceable half-space. Figure 1 intuitively shows the relationship between these points, half-spaces and the target set in a two-dimensional plane.
It should be noted that the half-space $\mathcal{H}_{t+1}^{\text{P}}$ can only prove that the distance from the average regret value $\bar{R}_{t}^{i}$ to the point $\boldsymbol{\psi}_{t}$ converges to 0, but $\boldsymbol{\psi}_{t}$ is not necessarily on the target cone $S^{i}$. So it cannot be shown that $\bar{R}_{t}^{i}$ will converge to $S^{i}$. The projection of $\bar{R}_{t}^{i}$ to the target cone $S^{i}$ is point $\boldsymbol{\gamma}_{t}^{i}=\left[\bar{R}_{t}^{i}\right]^{-},\bar{R}_{t}^{i,-}=\min\{0,\bar{R}_{t}^{i}\}$. We need to keep the distance from the $\bar{R}_{t}^{i}$ to point $\boldsymbol{\psi}_{t}$ and point $\boldsymbol{\gamma}_{t}^{i}$ within a certain range, and preference degree $\delta(a)$ just plays this role.
Define $a^{BR}=\arg\max_{a\in\mathcal{A}^{i}}\bar{R}^{i}_{t}(a)$, $a^{P}=\arg\max_{a\in\mathcal{A}^{i}}\delta(a)\bar{R}^{i}_{t}(a)$. The distance from the average regret value $\bar{R}^{i}_{t}$ to $\boldsymbol{\gamma}_{t}^{i}$ is $\|\bar{R}^{i,+}_{t}\|_{2}$, and the distance from $\bar{R}^{i}_{t}$ to the point $\boldsymbol{\psi}_{t}$ is $\delta^{i}(a^{P})\bar{R}^{i}_{t}(a^{P})$. Define $\text{dist}(x,y)$ as the distance between points $x$ and $y$.
$\text{dist}(\bar{R}^{i}_{t},\boldsymbol{\gamma}_{t})=\|\bar{R}^{i,+}_{t}\|_{2}\leq\sqrt{|\mathcal{A}^{i}|}\bar{R}^{i,+}_{t}(a^{BR}),$ (17)
at the same time,
$\text{dist}(\bar{R}^{i}_{t},\boldsymbol{\psi}_{t})=\delta^{i}(a^{P})\bar{R}^{i,+}_{t}(a^{P})\geq\delta^{i}(a^{BR})\bar{R}^{i,+}_{t}(a^{BR}),$ (18)
since $\delta^{i}(a^{BR})\geq 1$, so:
$\text{dist}(\bar{R}^{i}_{t},\boldsymbol{\gamma}_{t})\leq\frac{\sqrt{|\mathcal{A}^{i}|}\delta^{i}(a^{P})}{\delta^{i}(a^{BR})}\text{dist}(\bar{R}^{i}_{t},\boldsymbol{\psi}_{t}).$ (19)
Define $\delta^{*}=\max_{a\in\mathcal{A}^{i}}(\delta^{i}(a))$. Therefore, from Blackwell’s approachability, we know that $\text{dist}(\bar{R}^{i}_{t},\boldsymbol{\psi}_{t})\leq\frac{L\sqrt{|\mathcal{A}^{i}|}\delta^{*}}{\sqrt{t}}$. Finally, we get:
$\text{dist}(\bar{R}^{i}_{t},\boldsymbol{\gamma}_{t})=\frac{L{|\mathcal{A}^{i}|}\delta^{*}}{\sqrt{t}}.$ (20)
The convergence speed will linearly slow down as $\delta^{*}$ increases. As long as $\delta^{*}$ is set within a reasonable range, it will not significantly reduce the convergence speed of Pref-CFR.
4.3 Vulnerability CFR Will Converge to an $\epsilon$-NE
Figure 2: The differences in the selected forcing actions and forcing half-spaces of RM and $\beta$ vulnerability RM in the two-dimensional plane.
The proof for the Vulnerability CFR algorithm is more straightforward. In the original scenario, the historical strategy $\bar{\sigma}^{i}_{t}$ converging to a NE is equivalent to the regret $\bar{R}^{i}_{t}$ converging to the convex target cone $S^{i}=\mathbb{R}_{\leq 0}^{|\mathcal{A}^{i}|}$. This holds true because, given the opponent’s strategy $\bar{\sigma}^{-i}_{t}$, player $i$ identifies a strategy $\bar{\sigma}^{i}_{t}$ such that regardless of which pure strategy player $i$ selects, their payoff will not increase (since $\bar{R}^{i}_{t}(a)\leq 0$ for all $a\in\mathcal{A}^{i}$ as $t\rightarrow\infty$), thus satisfying the conditions for NE.
However, what if our target cone is not $S^{i}=\mathbb{R}_{\leq 0}^{|\mathcal{A}^{i}|}$, but instead $S^{i}=\mathbb{R}_{\leq\beta}^{|\mathcal{A}^{i}|}$ for some $\beta\geq 0$? This means that the regret of any pure strategy will not exceed $\beta$, indicating that we have identified a $\beta$-NE.
In the iteration, we only need to adjust the translation. The original projection point is $\boldsymbol{\gamma}_{t}=\bar{R}^{i,-}_{t}$, and the distance that needs to be reduced is given by:
$\text{dist}(\bar{R}^{i}_{t},\boldsymbol{\gamma}_{t})=\|\bar{R}^{i}_{t}-\boldsymbol{\gamma}_{t}\|_{2}=\|\bar{R}^{i,+}_{t}\|_{2}.$ (21)
Now, the projection point becomes $\boldsymbol{\gamma}_{t}^{\beta}=\bar{R}^{i,-\beta}_{t}$, where $\bar{R}^{i,-\beta}_{t}=\min\{\bar{R}^{i}_{t},\beta\}$. At this point, the distance that needs to be reduced is:
$\text{dist}(\bar{R}^{i}_{t},\boldsymbol{\gamma}_{t}^{\beta})=\|\bar{R}^{i}_{t}-\boldsymbol{\gamma}_{t}^{\beta}\|_{2}=\|\bar{R}^{i,+}_{t}-\beta\|_{2}.$ (22)
Thus, in the CFR iteration, we simply need to replace $\bar{R}_{t}^{i}(a)$ with $\bar{B}_{t}^{i}(a)=\bar{R}_{t}^{i}(a)-\beta$ to ensure that the final strategy converges to a $\beta$-NE. Figure 2 intuitively illustrates the relationship between these points, the half-spaces, and the target set in a two-dimensional plane.
Strictly speaking, in many of our subsequent experiments, both Pref-CFR and Vulnerability CFR are used simultaneously. However, Pref-CFR is the more critical component of the algorithm. Using Vulnerability CFR alone does not produce meaningful results. Therefore, even though Vulnerability CFR is utilized, it will not be mentioned in the name of the algorithm.
4.4 Analysis of Pref-CFR Algorithm
Pre-CFR offers a method for converging to different strategies. Nevertheless, during actual training, the settings for $\sigma(I)$ and $\beta$ still require careful configuration by experts.
  1. 1.
    Setting larger values for $\delta(I,a)$ and $\beta(I)$ can enhance the distinctiveness of the final strategy’s style. However, this may result in slower convergence speeds and a strategy that deviates significantly from the NE, making it susceptible to astute opponents.
  2. 2.
    While we can adjust $\delta(I,a)$ and $\beta(I)$ at the micro level, translating macro-level style characteristics recognized by the public into parameter settings for each information set is challenging.
To address the first problem, setting appropriate parameters for $\delta(I,a)$ and $\beta(I)$ can help the final strategy converge to a reasonable interval. For the parameter $\delta(I,a)$, experiments indicate that when $\delta(I,a)\leq 5$, the convergence speed remains acceptable, significantly increasing the likelihood of the preferred action being adopted in the final output strategy. Regarding $\beta(I)$, it can be determined based on the specifics of different games and the experience of human experts. For instance, an error margin of 25 mbb/h in Texas Hold’em is typically inconspicuous.
The second problem is relatively more complex. Fortunately, for common games like Texas Hold’em, there is ample expert knowledge to draw upon. If a player has a narrow range of hands when entering the pot, their style is classified as "tight"; conversely, a wider range is termed "loose." For definitions of these Texas Hold’em terms, please refer to Appendix LABEL:sec:game-introduction. If a player exhibits a relatively high proportion of 3-bets after entering the pot, they are labeled as "aggressive." These styles exhibit strong consistency. Thus, we can apply the same set of $\delta$ values across all information sets. For example, to make the AI more aggressive, we could increase $\delta(I,\text{Raise})$ for all raising actions across all information sets.
However, for more extensive-form games that lack extensive expert analysis, the design of $\delta(I,a)$ and $\beta(I)$ remains an area requiring further research.
5 Experiments
Our experiments are conducted using Kuhn poker Kuhn (1950) as well as two-player and three-player Texas Hold’em Poker. A brief introduction to the rules of these games can be found in Appendix LABEL:sec:game-introduction. The Kuhn poker experiment has been made publicly available on GitHub.
In our Kuhn poker experiments, we used the vanilla Pref-CFR while Texas Hold’em experiments utilized a variant of the multi-valued state technique Brown et al. (2018). Solutions were computed in under 10 minutes with a 24-core CPU; subgames included 35k states and the full game used for leaf estimates had 25M states. In Heads-Up play, this setup achieved exploitability below 4 mBB/h. For reference, Brown et al. Brown et al. (2018) found that 30 seconds on a 4-core CPU met their AI’s requirements.
5.1 Pref-CFR converges to different equilibria
Before analyzing the Pref-CFR algorithm, it is essential to first illustrate the situations in which CFR fails to converge to different NEs. We will use Kuhn poker as an example. In this game, player 1 faces multiple equilibria, while player 2 has only one equilibrium strategy. The equilibrium strategy for player 1 in Kuhn poker is detailed in Table 1.
Table 1: The equilibrium strategy of player 1 in Kuhn poker, where $\alpha\in[0,1/3]$. It can be considered that player 1 has countless equilibrium strategies in this game.
Here’s a more concise version of your paragraph:
[…]
Figure 3: Convergence rate of CFR in Kuhn poker (left) and fluctuation of $\alpha$ in CFR algorithm iterations (right). A total of 100 experiments were conducted, and all experiments were carried out with a fixed 10,000 iterations. The shaded areas represent the 90% confidence intervals for these trials. It can be seen that all CFR iterations will converge to $\alpha=0.2$.
We generate the initial strategy for CFR training using a random distribution. As shown in Figure 3, while Kuhn poker has numerous equilibria, CFR appears to converge solely to one equilibrium at $\alpha=0.2$, regardless of the initial strategy. This phenomenon is not unique to Kuhn poker; it seems to be a common issue for CFR and its variants. Moreover, the strategy produced by the CFR algorithm aligns closely with the traditional "machine strategy," characterized by complexity and stability without distinct features. In real games, human players can easily identify and remember simpler strategies, such as $\alpha=0$ and $\alpha=1/3$. For instance, when $\alpha=0$, player 1 consistently chooses Pass, projecting a cautious and conservative image. We aim for Pref-CFR to learn such strategies, prompting us to design six sets of parameters for it.
  1. 1.
[…]
  2. 2.
[…]
  3. 3.
[…]
  4. 4.
[…]
  5. 5.
[…]
  6. 6.
[…]
Figure 4: Convergence rate of CFR/Pref-CFR in Kuhn poker (left) and fluctuation of $\alpha$ value in CFR/Pref-CFR iterations (right). Thirty experiments were conducted for each setting. All experiments were carried out with a fixed 10,000 iterations. The shaded areas represent the 90% confidence intervals for these trials. It can be seen that the Pref-CFR algorithm can still converge to equilibrium, and the convergence speed is comparable to that of the original CFR. In addition, the right figure clearly shows that with the change of the parameter design of Pref-CFR, the final strategy successfully converges to different NEs.
Except for the preference degree settings mentioned above, the preference degrees on all other information sets are $1$. Running 10,000 iterations, the final results are displayed in Figure 4. In terms of convergence speed, all strategies, except for the most aggressive setting (1), exhibit convergence rates similar to the original CFR. Regarding equilibrium convergence, all Pref-CFR variations yield equilibria that deviate from $\alpha=0.2$, with greater preference degree settings resulting in more pronounced deviations. Additionally, the performance of Pref-CFR(RM) is considerably inferior to that of Pref-CFR(BR). Thus, we prefer using the Pref-CFR(BR) method.
5.2 Pref-CFR converges to different styles in Texas Hold’em
Although a NE strategy can yield the best reward in a worst-case scenario, it often does not align with people’s expectations. Consider the following choices: first, to receive \$4,500 with 100% certainty; second, to have a 50% chance of winning \$10,000, but also a 50% chance of winning nothing. While the expected payoff of the second option is higher, most people prefer the guaranteed \$4,500. This preference stems from the fact that, compared to the potential risk, losing \$500 is within a tolerable range.
The traditional CFR algorithm does not account for goals beyond expected payoffs. In Texas Hold’em, two significant "style demands" have consistently emerged, yet the previous CFR algorithm was unable to address them. Consequently, we have designed two strategies for Texas Hold’em training:
  1. 1.
    Aggressive strategy A prominent feature of this strategy is its higher probability of raising. For some amateur players with an abundance of chips, sensitivity to chip loss is minimal. Their primary motivation for participating in Texas Hold’em is not necessarily to win more chips but rather to seek novel experiences or enhance the atmosphere at social gatherings. As a result, these players often look forward to engaging in games with larger pot amounts. In training, we set $\delta(I,raise)=5$ and $\beta=0.05$ at the first decision node of player 1.
  2. 2.
    Loose passive strategy A significant characteristic of this strategy is that the probability of folding is lower than that of a typical strategy. This phenomenon arises from the difficulty many players with a small number of chips experience when it comes to folding. After investing a considerable amount of chips during the flop or turn stages, these players often feel uncomfortable quitting the game without knowing their opponent’s hole cards. The discomfort is especially pronounced when they suspect they have been successfully bluffed. By minimizing folds, this strategy provides psychological comfort for these players, helping them avoid the frustration of being outplayed by an opponent’s bluff. In training, we set $\delta(I,call)=5$ and $\beta=0.05$ at the first decision node of player 1.
Table 2: The battle results between different AIs in two-player Texas Hold’em. In Texas Hold’em, mBB/h represents the thousandth of a big blind won or lost per hand, used to accurately measure a player’s profit or loss for each individual hand.
Figure 5: Strategy display for two-player Texas Hold’em. In the top left corner of each image, the current player’s information and available actions are displayed. The central area showcases the strategies for different hand combinations at this stage. In Texas Hold’em poker, there are 13 ranks across 4 suits, with no distinction in value between suits, resulting in 169 unique hand combinations. These are represented in a 13$\times$13 matrix, where the lower left displays offsuit hands and the upper right shows suited hands. Each matrix element’s color indicates the strategic choice for the corresponding hand: blue for folding, green for calling, red shades for raising (with deeper red shades indicating higher raises), and black-red for going all-in. The bottom row provides an overview of the average strategies across all hands, allowing for a visual understanding of the overall strategy distribution.
Before analyzing the style variations among different AI models, we first examine the performance differences between them. We aim to develop AIs that embody distinct play styles without sacrificing significant winnings. Given the challenge of evaluating strategy effectiveness in large-scale games, we use head-to-head matches to measure the capabilities of different algorithms. The results, shown in Table 2, reveal that in both two-player and three-player games, the discrepancies in performance between different styles are minimal, with differences within 10mBB/h. Previous research has established that a training error under 1mBB/h indicates convergence to a NE Bowling et al. (2015). For context, in the renowned matches between Pluribus and top human players, Pluribus achieved an average win rate of 32mBB/h Brown and Sandholm (2019b). Thus, the 10mBB/h difference observed in our strategies suggests that these AIs possess equivalent skill levels.
Our experiments clearly demonstrate how the algorithm influences the final strategies in Figure 5. In standard CFR training for two-player Texas Hold’em, the average preflop strategy distribution is [5.4%, 52.7%, 42.0%, 0.0%, 0.0%] for folding, calling, raising to 2, raising to 3, and going all-in, respectively. In contrast, the loose AI’s strategy distribution shifts to [0.3%, 73.3%, 21.7%, 4.7%, 0.0%], while the aggressive AI shows [4.2%, 64.8%, 9.8%, 21.1%, 0.0%]. As expected, the loose AI exhibits a much lower folding probability, dropping from 5.4% to 0.3%—a 94.4% decrease. The differences between aggressive and standard strategies are also pronounced, with the aggressive AI raising to 3 at a 21.1% rate compared to 0% in the standard strategy.
We believe that these AI strategies offer valuable insights for human players. For example, traditional human experts and GTO-based AIs typically fold weak hand combinations like 82o and 72o, which are considered poor due to their low potential for straights, lack of suitedness, and absence of high cards. However, our loose-passive AI opts to call almost 100% of the time, and the aggressive AI may even raise with these hands. Such stylized strategies cleverly adjust hand distributions to obscure the strength of the hands, akin to "hiding a leaf in the forest." The loose-passive AI’s increased calling rates make it challenging for opponents to discern whether a weak hand is being bluffed or a strong hand is being slow-played. Similarly, the aggressive AI increases the raising frequency across all hand combinations, potentially altering the conventional approach to Texas Hold’em. While the loose-passive style was historically viewed as suboptimal, our findings suggest that, when balanced across all hand combinations, it can be a viable strategy, offering a fresh perspective on game tactics.
Table 3: The battle results between different AIs in three-player Texas Hold’em.
Figure 6: Strategy display for three-player Texas Hold’em. The understanding of these strategies is the same as that for two-player games and will not be repeated.
The results of three-player AI matches, displayed in Table 3 and Figure 6, show that our algorithm achieves even more pronounced effects in three-player scenarios. In standard three-player training, the average strategy distribution is [61.4%, 0.0%, 38.6%, 0%]. This example illustrates why the loose-passive style was previously considered ineffective: according to GTO calculations, it is typically not recommended. However, our experiments reveal that the Loose-Passive strategy not only avoids any loss in earnings but actually gains an advantage of 0.2mBB/h compared to the Normal strategy, while significantly increasing the calling probability from 0.0% to 23.9%. We speculate that the added complexity of three-player poker compared to two-player games accentuates the impact of style changes.
6 Conclusion and Prospect
The Pref-CFR algorithm introduced in this paper addresses the limitations of the previous CFR algorithm in achieving diverse equilibria. Our experiments successfully trained AIs with distinct playing styles for Texas Hold’em. For instance, the aggressive AI increased its probability of raising to 3 from 0.0% to 16.0% in three-player Texas Hold’em, with only a minimal loss of 10mBB/h. Meanwhile, the loose AI increased its calling probability from 0.0% to 23.9%. Additionally, our algorithm does not significantly slow down training and can provide real-time strategic guidance for human players during gameplay.
On the other hand, the current approach requires manually setting preference degrees for actions in each information set, limiting the algorithm’s flexibility in accommodating different player styles. Given the game’s complexity, it is challenging for players to adopt different strategies across various situations. Enhancing the algorithm’s practicality would involve developing a method to automatically convert user-defined metrics (such as betting ratios or pot entry rates) into preference degrees for each information set, thereby streamlining the customization process.
Appendix A GAME INTRODUCTION
A.1 Kuhn poker
Kuhn poker is a simple poker game. Here is a detailed rules introduction to it:
  • Card composition: Kuhn poker uses only three cards: J, Q, and K. Each player can only get one card at a time.
  • Action sequence: Usually, two players participate in the game. At the beginning of each round, each player places a blind bet first. Then each player will randomly receive a card. Subsequently, players take actions in turn. The action options are Bet and Pass. If one player bets, the other player can choose to call(Bet) or fold(Pass).
  • Winning determination: Finally, if a player calls, both players show their cards and compare the sizes. The player with the larger card wins and takes the chips in the pot. In Kuhn poker, K > Q > J.
Kuhn poker is a classic model in game theory research. Due to its simple rules and limited card types and action choices, it is convenient for mathematical analysis and theoretical derivation.
A.2 Texas Hold’em
Texas Hold’em is a poker game that has a large player population and extensive influence globally. In terms of tournaments, numerous international Texas Hold’em competitions draw significant attention. High prize money attracts many top players to participate. Socially, it is a popular activity for people to enhance communication during leisure gatherings. The following is an introduction to some basic Texas Hold’em poker terminology. For more content, you can refer to Wikipedia.
  • Preflop: The stage before the community cards are dealt. Decisions are made based on hands, position, etc., like raising, calling or folding.
  • Flop: The first three community cards dealt. Evaluate competitiveness combined with hands and decide strategies.
  • Turn: The fourth community card after the flop.
  • River: The last community card. Decides the final decision.
  • Big blind (BB): The big blind is a forced bet made by one of the players before the cards are dealt. It is typically twice the size of the small blind.
  • Small blind (SB): The small blind is also a forced bet made by a player before the cards are dealt.
  • Button: The button indicates which player is the dealer for that hand. The player on the button has certain advantages in terms of position and play order.
  • mBB/h: In Texas Hold’em, mBB/h represents the thousandth of a big blind won or lost per hand, used to accurately measure a player’s profit or loss for each individual hand.
  • Loose: Play many hands with a low entry criterion. Often participate even with weak hands. High risk but may win with weak hands against strong ones.
  • Aggressive: Actively bet or raise to put pressure on opponents and control the game to win the pot. The hand may not be strong.
  • 3 Bet: Raise again after someone has raised. Often indicates a strong hand or wanting to pressure opponents and increase the pot.
  • All-in: Bet all chips. Due to confidence in the hand or having few chips and wanting to pressure.
  • Call: Follow by betting the same amount of chips as the opponent. Want to continue to see the cards and compete.
  • Fold: Give up the hand and not participate in the pot. Due to weak hands or high risk from large bets by opponents.
  • Bluff: When having a weak hand, make large bets and let opponents think the hand is strong so they fold to win the pot. High risk.